OpenStudy (anonymous):

simplify the expression. write the answer in exponent form. t^7xt^7

OpenStudy (calculusxy):

Do you mean: \(\large t^7 \times t^7\)

OpenStudy (calculusxy):

If you do mean that then since this expression has two bases that are the same you just need to add the exponents. \(\large \color {purple} {x^a \times x^b = x^{a + b}}\)